OpenStudy (anonymous):

For the following system, if you isolated y in the first equation to use the Substitution Method, what expression would you substitute into the second equation? 4x − y = 8 2x + 3y = 9

OpenStudy (anonymous):

isolating Y on the first one you get: -y = 8 - 4x y = -8 + 4x ------ (equation 3) so if you that Y is equal that. Substitute equation 3 into the second one (the y value): 2x + 3(-8 + 4x) = 9 2x -24 + 12x = 9 14x = 9 + 24 14x = 33 x = 33/14 am i right?

hero (hero):

y = 4x - 8 y = (9-2x)/3 y = y 3(4x-8) = 9 - 2x 12x - 24 = 9 - 2x 14x = 33 x = 33/14